# Five-loop Anomalous Dimensions of Cubic Scalar Theory from Operator Product Expansion

### v2

@@ 4 Five-loop anomalous dimensions
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\bigg{(}\frac{7551}{8}-702\zeta_{3}^{2}-2322\zeta_{3}-1215\zeta_{5}+\frac{5\pi^{6}}{7}+\frac{72\pi^{4}}{5}\bigg{)}Q^{3}\bigg{]}\epsilon^{5}~,$
and expressions $\Delta_{Q{\tiny\mbox{FP}}}^{4},\Delta_{Q{\tiny\mbox{FP}}}^{5}$ are presented in Appendix §C. As a confirmation of these results, the 3-loop contribution up to $\epsilon^{3}$ order is consistent with literature findings Vasiliev:1981yc ; Vasiliev:1981dg ; Vasiliev:1982dc ; Jack:2021ziq  [^7] The value $a_{t}^{(3)}$ in paper Jack:2021ziq that computes $\Delta_{Q}^{(3)}$ is incorrect. After using the correct value $a_{t}^{(3)}=\frac{g^{6}}{3}+g^{5}h$, the $\Delta_{Q}^{(3)}$ and the $1/N^{3}$ term in their paper are consistent with the 3-loop order part of our results.. The corresponding terms $\frac{Q^{i+1}\epsilon^{i}}{N^{i}},\frac{Q^{i+1}\epsilon^{i+1}}{N^{i}},i=1,\ldots,5$ are also consistent with those reported in the semi-classical literature Antipin:2021jiw . The 4-loop contribution up to $1/N^{2}$ order has been previously provided in Huang:2024hsn , while the 5-loop contributions up to $1/N^{5}$ order presented here are new. Our results have been provided in ancillary files to the arXiv submission of this article.
@@ 5 Conclusion
In this paper, we apply the OPE method of renormalization proposed in Huang:2024hsn to evaluate the 5-loop scaling dimensions of the $\phi^{Q}$ operator in the six-dimensional cubic scalar theory. The result is derived from over 40 000 two-point propagator-type integrals computed using the graphical function method, combined with non-trivial summation among these integrals. We have successfully obtained the 5-loop correction to the anomalous dimension of the $\phi^{Q}$ operator, as well as the large $N$ expansion of the scaling dimension at the Wilson-Fisher fixed point up to the $1/N^{5}$ order. It stands for a new record for the scaling dimensions of $\phi^{Q}$ operator in cubic scalar theory. The result shows consistency with lower-loop perturbative calculations reported in other literature and aligns with the corresponding terms derived via semi-classical methods. Furthermore, the non-trivial cancellation of $\epsilon$ poles in the final result provides a strong consistency check validating its correctness. This work further demonstrates the efficiency and power of the OPE method in renormalization computations.
@@
Efforts to apply the OPE method to address the challenges posed by more complex problems are currently underway. A promising project aims to surpass the 6-loop record in quartic scalar theory by computing the 7-loop scaling dimensions of the $\phi^{Q}$ operator. Such research would not only advance the precision of renormalization computations but also drive the evolution of modern techniques for multi-loop integral evaluations. Also as previously noted, the OPE method is highly general and extendable beyond scalar theories. It is interesting to investigate the renormalization computations by OPE method in other situations, such as the Gross-Neveu-Yukawa theory with fermions, or scalar-QED theory with gauge fields, etc. In these theories, Feynman integrals are no longer purely scalar integrals, but involve non-trivial numerators. While further developments may enable the computation of such non-trivial master integrals via graphical function method adapted to numerator structures, the most viable current approach relies on the computational algorithms Smirnov:2008iw ; Maierhofer:2017gsa ; Wu:2023upw ; Guan:2024byi based on the IBP method Tkachov:1981wb ; Chetyrkin:1981qh ; Laporta:2000dsw . Another potential challenge would be applying the OPE method to the anomalous dimensions of more intricate operators, such as those with higher derivatives, Lorentz indices, or in scenarios with operator mixing Derkachov:1997ch ; Cao:2021cdt . While the application of the OPE method remains straightforward in these cases Huang:2024hsn , practical computations demand more sophisticated techniques. It is interesting to explore the diverse applications of the OPE method in renormalization, and further enhance its capabilities.
@@ Acknowledgments
We would like to thank Oliver Schnetz for stimulating conversations on the graphical function method and the usage of HyperlogProcedures package. This work is supported by the Science Challenge Project (No. TZ2025012), and NSAF No.U2330401.
@@ Appendix C Large $N$ expansion at the Wilson-Fisher fixed point
Here we present the results of the $1/N^{4}$ and $1/N^{5}$ orders in the large $N$ expansion of $\gamma_{Q}$ at Wilson-Fisher fixed point. We have,

### v4

@@ 4 Five-loop anomalous dimensions
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\bigg(\frac{7551}{8}-702\zeta_{3}^{2}-2322\zeta_{3}-1215\zeta_{5}+\frac{5\pi^{6}}{7}+\frac{72\pi^{4}}{5}\bigg)Q^{3}\bigg]\epsilon^{5}~.$
Expressions $\Delta_{Q{\tiny\mbox{FP}}}^{4}$ and $\Delta_{Q{\tiny\mbox{FP}}}^{5}$ are presented in Appendix §C. These results are further validated by the following consistency checks. The 3-loop contribution up to $\epsilon^{3}$ order is consistent with literature findings Vasiliev:1981yc ; Vasiliev:1981dg ; Vasiliev:1982dc ; Jack:2021ziq  [^8] The value $a_{t}^{(3)}$ in paper Jack:2021ziq that computes $\Delta_{Q}^{(3)}$ is incorrect. After using the correct value $a_{t}^{(3)}=\frac{g^{6}}{3}+g^{5}h$, the $\Delta_{Q}^{(3)}$ and the $1/N^{3}$ term in their paper are consistent with the 3-loop order part of our results.. The corresponding terms $\frac{Q^{i+1}\epsilon^{i}}{N^{i}},\frac{Q^{i+1}\epsilon^{i+1}}{N^{i}},i=1,\ldots,5$ are also consistent with those reported in the semi-classical literature Antipin:2021jiw . The 4-loop contribution up to $1/N^{2}$ order has been previously provided in Huang:2024hsn , while the 5-loop contributions up to $1/N^{5}$ order presented here are new. All the above results have been provided in ancillary files to the arXiv submission of this article.
@@ 5 Conclusion
In this paper, we employ the OPE based renormalization algorithm proposed in Huang:2024hsn to compute the 5-loop scaling dimensions of the $\phi^{Q}$ operator in the six-dimensional cubic scalar theory. This algorithm represents a non-trivial implementation of the connection between Wilson coefficients of OPE and anomalous dimensions, enabling the determination of renormalization $Z$-factors by treating UV divergences globally. A key advantage is its avoidance of cumbersome sub-divergence subtraction. Instead, it systematically extracts the essential UV divergences for renormalization directly from two-point integrals. Advancements in modern computational techniques, particularly the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP adopted herein, have rendered feasible the evaluation of these two-point integrals even beyond five loops. The transformative impact of the graphical function method, combined with the efficiency and robustness of the proposed OPE based algorithm, has facilitated the successful completion of renormalization computations up to five loops on a standard desktop computer. These results are derived from over 40,000 two-point propagator-type integrals computed using the HyperlogProcedures package HP , augmented by non-trivial summation protocols inherent to the algorithm. We have successfully obtained the 5-loop corrections to the anomalous dimension of the $\phi^{Q}$ operator, as well as the large $N$ expansion of the scaling dimension at the Wilson-Fisher fixed point up to $1/N^{5}$ order. It stands for a new record for the scaling dimensions of the $\phi^{Q}$ operator in cubic scalar theory. Our results exhibit perfect consistency with lower-loop perturbative calculations reported in the literature and align with corresponding terms derived via semi-classical methods. Furthermore, the non-trivial cancellation of $\epsilon$ poles in the final result provides a stringent consistency check, validating its correctness. This work further demonstrates the efficiency and versatility of the OPE based algorithm for high-precision renormalization computations in quantum field theories.
@@
Efforts are currently underway to apply this OPE based algorithm to tackle the challenges presented by more sophisticated problems. One promising initiative aims to break the existing 6-loop benchmark in quartic scalar theory by computing the 7-loop scaling dimensions of the $\phi^{Q}$ operator. Such research will not only push the precision of renormalization computations but also drive the evolution of modern techniques for multi-loop integral evaluations. Also as previously noted, the OPE based algorithm boasts strong generality and extensibility beyond scalar field theories. It would be particularly insightful to explore its application to renormalization computations in other theoretical frameworks, such as the Gross-Neveu-Yukawa theory and scalar QED, among others. In these theories, Feynman integrals are no longer purely scalar in nature, but instead involve non-trivial numerator structures. While future methodological refinements may enable the calculation of such non-trivial master integrals via a modified graphical function method tailored to handle numerator-dependent topologies, the most robust approach available at present relies on computational algorithms Smirnov:2008iw ; Maierhofer:2017gsa ; Wu:2023upw ; Guan:2024byi based on the integration-by-parts (IBP) method Tkachov:1981wb ; Chetyrkin:1981qh ; Laporta:2000dsw . Another potential challenge lies in extending this algorithm to compute the anomalous dimensions of more intricate operators, for instance, those containing higher derivatives, Lorentz indices, or in scenarios involving operator mixing Derkachov:1997ch ; Cao:2021cdt . Although the conceptual application of the OPE based algorithm remains straightforward in these cases, as demonstrated in Huang:2024hsn , practical computations necessitate more advanced technical implementations. Further exploration of the algorithm’s diverse applications in renormalization and targeted enhancements of its computational capabilities will undoubtedly open new avenues for high-precision quantum field theory calculations.
@@ Acknowledgments
We would like to thank Oliver Schnetz for stimulating conversations on the graphical function method and the usage of HyperlogProcedures package, and Hugh Osborn for useful comments on this paper. This work is supported by the Science Challenge Project (No. TZ2025012), and NSAF No.U2330401.
@@ Appendix C Large $N$ expansion at the Wilson-Fisher fixed point
Here we present the results of the $1/N^{4}$ and $1/N^{5}$ orders in the large $N$ expansion of $\gamma_{\varphi^{Q}}$ at Wilson-Fisher fixed point. We have,